# Bayesian Hierarchical Models for Quantitative Estimates for Performance metrics applied to Saddle Search Algorithms

### v2

@@ Abstract
Rigorous performance evaluation is essential for developing robust algorithms for high-throughput computational chemistry. Traditional benchmarking, however, often struggles to account for system-specific variability, making it difficult to form actionable conclusions. We present a Bayesian hierarchical modeling framework that rigorously quantifies performance metrics and their uncertainty, enabling a nuanced comparison of algorithmic strategies. We apply this framework to analyze the Dimer method, comparing Conjugate Gradient (CG) and L-BFGS rotation optimizers, with and without the removal of external rotations, across a benchmark of 500 molecular systems. Our analysis confirms that CG offers higher overall robustness than L-BFGS in this context. While the theoretically-motivated removal of external rotations led to higher computational cost (¿40
@@ 1 Introduction
Locating transition states (TS), first-order saddle points on potential energy surfaces (PES), is fundamental to understanding reaction mechanisms and kinetics in chemistry [10, 40]. Numerous algorithms target this challenge, with minimum mode following (MMF) methods forming a prominent class [7]. These algorithms, initiated from a single point on the PES, iteratively ascend the minimum curvature mode while relaxing orthogonal directions. Notable MMF families include Lanczos-based Activation Relaxation Technique Nouveau (ARTn) methods [34, 5, 35, 41, 18] and the two-point Dimer method [19], which uses a pair of images to estimate the minimum mode direction. Other variations employ internal coordinates [20] or iteratively build PES approximations [26].
@@ 2.1 Dimer method
The Dimer method [19] is a single-ended saddle search method that algorithmically finds first-order saddle point geometries by iteratively refining the position $R$ (a 3N-dimensional vector representing the system coordinates) and orientation $\hat{N}$ (a unit vector in $3N$ dimensions) of a “dimer”. The dimer consists of two system images, $R_{1}$ and $R_{2}$, separated by a small fixed distance $\Delta R$ along the dimer axis $\hat{N}$:
@@ 3 Applications
To demonstrate the utility of leveraging a Bayesian approach to performance statistices, we analyse the performance of the dimer method variants on the dataset of 500 initial configurations for small gas-phase organic molecules (7-25 atoms) introduced by hermesSellaOpenSourceAutomationFriendly2022 [20]. Saddle point searches were conducted using the EON software package [9]  [^1] Using the Github variant: https://github.com/TheochemUI/eOn, interfaced with NWChem [1] as the quantum mechanical engine. Energies and forces were calculated at the Hartree-Fock (HF) level of theory using RHF (restricted HF) for singlets and UHF (unrestricted HF) for doublets [24] with the 3-21G basis set. Saddle searches algorithms were considered converged when the maximum component of the atomic forces fell below 0.01 eV/Å. Comprehensive details regarding computational parameters, including self-consistent field thresholds, specific EON settings, and interface configurations, are provided in the Supplementary Information.
@@ 3.1 Data overview
Figures 2 and 3 illustrate the difference in PES calls required for convergence, comparing pairs of calculations that reached the same saddle point (within 0.01 eV). When rotation removal was disabled (Fig. 2, 440 comparable systems), the CG-rotations dimer required fewer PES calls than L-BFGS-rotations in a majority of cases (55%, median reduction 21 calls), although L-BFGS is rather comparable in performance (in 43% cases, median reduction 16 calls). With CG-rotations (Fig. 3, 465 comparable systems), disabling rotation removal was markedly more efficient, requiring fewer PES calls than enabling it in over 99% of systems (median reduction 126 calls). Both figures indicate that performance differences tend to be larger for systems starting with higher initial RMSD values relative to the saddle point.
@@ 3.2 Effort analysis model
$\eta_{ij}$ =$=\begin{cases}\beta_{0}+\beta_{1}\text{DR}_{i(j))}+u_{j}&\text{(a)}\\ \beta_{0}+\beta_{2}\text{RR}_{i(j))}+u_{j}&\text{(b)}\\ \beta_{0}+\beta_{1}\text{DR}_{i(j))}+\beta_{2}\text{RR}_{i(j))}+\beta_{3}(\text{DR}_{i(j))}\times\text{RR}_{i(j))})+u_{j}&\text{(c)}\end{cases}$
@@ Acknowledgments
The full computational workflow, analysis code, and data supporting this study are openly available. A version-tagged GitHub repository (https://github.com/HaoZeke/brms_idrot_repro) provides all resources needed to reproduce the findings. To ensure full reproducibility, the repository contains a complete computational environment managed by pixi, with version-pinned software dependencies including EON, NWChem, and ASE, alongside utilities from Wailord [13]. In line with best practices for literate and reproducible programming [16, 17, 42, 15], some dependencies are bundled directly via gitsubrepo. For permanent access and citation, the final curated dataset is hosted on the Materials Cloud Archive [44] under the entry goswamiBayesianHierarchicalModels2025MCDat [14].

### v3

@@ Abstract
Rigorous performance evaluation is essential for developing robust algorithms for high-throughput computational chemistry. Traditional benchmarking, however, often struggles to account for system-specific variability, making it difficult to form actionable conclusions. We present a Bayesian hierarchical modeling framework that rigorously quantifies performance metrics and their uncertainty, enabling a nuanced comparison of algorithmic strategies. We apply this framework to analyze the Dimer method, comparing Conjugate Gradient (CG) and L-BFGS rotation optimizers, with and without the removal of external rotations, across a benchmark of 500 molecular systems. Our analysis confirms that CG offers higher overall robustness than L-BFGS in this context. While the theoretically-motivated removal of external rotations led to higher computational cost (¿40% more energy and force calls) for most systems in this set, our models also reveal a subtle interplay, hinting that this feature may improve the reliability of the L-BFGS optimizer. Rather than identifying a single superior method, our findings support the design of adaptive “chain of methods” workflows. This work showcases how a robust statistical paradigm can move beyond simple performance rankings to inform the intelligent, context-dependent application of computational chemistry methods.
@@ 1 Introduction
Locating transition states (TS), first-order saddle points on potential energy surfaces (PES), is fundamental to understanding reaction mechanisms and kinetics in chemistry [10, 42]. Numerous algorithms target this challenge, with minimum mode following (MMF) methods forming a prominent class [7]. These algorithms, initiated from a single point on the PES, iteratively ascend the minimum curvature mode while relaxing orthogonal directions. Notable MMF families include Lanczos-based Activation Relaxation Technique Nouveau (ARTn) methods [35, 5, 36, 43, 19] and the two-point Dimer method [20], which uses a pair of images to estimate the minimum mode direction. This is in turn an independent implementation of the hybrid eigenvector-following algorithm [37]. Other variations employ internal coordinates [21] or iteratively build PES approximations [27, 16].
@@ 2.1 Dimer method
The Dimer method [20] is a single-ended saddle search method that algorithmically finds first-order saddle point geometries by iteratively refining the position $R$ (a 3N-dimensional vector representing the system coordinates) and orientation $\hat{N}$ (a unit vector in $3N$ dimensions) of a “dimer”. An alternative formulation of the same was demonstrated earlier via the hybrid-eigenvector following using the variational Rayleigh-Ritz to find the eigenvalue and eigenvector pair [37]. The dimer consists of two system images, $R_{1}$ and $R_{2}$, separated by a small fixed distance $\Delta R$ along the dimer axis $\hat{N}$:
@@ 2.2.1 Effect structures considered
$\text{Fixed Effects}_{ij}=\begin{cases}\beta_{0}+\beta_{1}\text{DR}_{i(j))}\quad\text{(RotOptimizer)}\\ \beta_{0}+\beta_{2}\text{RR}_{i(j))}\quad\text{(RotRemoval)}\\ \beta_{0}+\beta_{1}\text{DR}_{i(j))}+\beta_{2}\text{RR}_{i(j))}\\ \quad+\beta_{3}(\text{DR}_{i(j))}\times\text{RR}_{i(j))})\quad\text{(Full)}\end{cases}$ (6)
@@ 3 Applications
To demonstrate the utility of leveraging a Bayesian approach to performance statistices, we analyse the performance of the dimer method variants on the dataset of 500 initial configurations for small gas-phase organic molecules (7-25 atoms) introduced by hermesSellaOpenSourceAutomationFriendly2022 [21]. Saddle point searches were conducted using the EON software package [9]  [^1] Using the Github variant: https://github.com/TheochemUI/eOn, interfaced with NWChem [1] as the quantum mechanical engine. Energies and forces were calculated at the Hartree-Fock (HF) level of theory using RHF (restricted HF) for singlets and UHF (unrestricted HF) for doublets [25] with the 3-21G basis set. Saddle searches algorithms were considered converged when the maximum component of the atomic forces fell below 0.01 eV/Å. This rather loose tolerance was chosen to highlight the effect of rotation removal, as the step sizes decrease on approaching the saddle and are less likely to lead to spurious rotation. Additionally, tighter tolerances tend to put more focus on the parameters of the underlying computational engine which is not the primary goal here. Comprehensive details regarding computational parameters, including self-consistent field thresholds, specific EON settings, and interface configurations, are provided in the Supplementary Information.
@@ 3.1 Data overview
Figures 1 and 2 illustrate the difference in PES calls required for convergence, comparing pairs of calculations that reached the same saddle point (within 0.01 eV). When rotation removal was disabled (Fig. 1, 440 comparable systems), the CG-rotations dimer required fewer PES calls than L-BFGS-rotations in a majority of cases (55%, median reduction 21 calls), although L-BFGS is rather comparable in performance (in 43% cases, median reduction 16 calls). With CG-rotations (Fig. 2, 465 comparable systems), disabling rotation removal was markedly more efficient, requiring fewer PES calls than enabling it in over 99% of systems (median reduction 126 calls). Both figures indicate that performance differences tend to be larger for systems starting with higher initial RMSD values relative to the saddle point.
@@
It is important to note that there are alternative methods for removing global rotations and translations, such as those implemented in OPTIM [47]. In OPTIM, the Rayleigh-Ritz procedure [37] and subsequent partial tangent space minimization involve orthogonalization to overall translation or rotation. However, the application demonstrated here follows the quaternion-based method implemented in EON [33].
@@ 3.2 Effort analysis model
$\eta_{ij}$ =$=\begin{cases}\beta_{0}+\beta_{1}\text{DR}_{i(j))}+u_{j}&\text{(a)}\\ \beta_{0}+\beta_{2}\text{RR}_{i(j))}+u_{j}&\text{(b)}\\ \beta_{0}+\beta_{1}\text{DR}_{i(j))}\\ +\beta_{2}\text{RR}_{i(j))}+\beta_{3}(\text{DR}_{i(j))}\times\text{RR}_{i(j))})+u_{j}&\text{(c)}\end{cases}$
@@ 4 Discussion
While this study has focused on variants of the Dimer method, the Bayesian hierarchical framework presented here is broadly applicable. An important future direction is its application to a comparative benchmark of fundamentally different classes of algorithms, such as single-ended methods (Dimer, Sella) versus chain-of-states methods (NEB, GSM). Such an analysis could quantitatively address long-standing questions, for example, by modeling the probability of finding the correct path as a function of the number of images in an NEB calculation or comparing the raw computational cost versus the likelihood of success for finding a transition state starting from only the reactant and product states. This would provide the community with a robust, data-driven basis for selecting the optimal algorithm class for a given high-throughput workflow.
@@ Acknowledgments
The full computational workflow, analysis code, and data supporting this study are openly available. A version-tagged GitHub repository (https://github.com/HaoZeke/brms_idrot_repro) provides all resources needed to reproduce the findings. To ensure full reproducibility, the repository contains a complete computational environment managed by pixi, with version-pinned software dependencies including EON, NWChem, and ASE, alongside utilities from Wailord [13]. In line with best practices for literate and reproducible programming [17, 18, 44, 15], some dependencies are bundled directly via gitsubrepo. For permanent access and citation, the final curated dataset is hosted on the Materials Cloud Archive [46] under the entry goswamiDatasetBayesianHierarchical2025 [14].